OpenStudy (lidia.12):

Find the inverse of this equation. Y=(x+5)/2 Can anyone help?

OpenStudy (mhchen):

To find the inverse, you flip x and y: x = (y+5)/2 Then solve for y: 2x = y + 5 And I think you can continue from here..

KjSaif (kjsaif):

\[f^-1 (x)=-5+2x\]

KjSaif (kjsaif):

@Lidia.12 it's \[f^-1 (x)=-5+2x\]

OpenStudy (lidia.12):

I'm still a bit lost